OpenStudy (anonymous):

Use the laws of exponents to simplify. (6^3/14)^2/11

OpenStudy (anonymous):

3/14 * 2/11 = 6/154 = 3/77 answer = 6 ^(3/77)

OpenStudy (anonymous):

ok ? - doesn't look very 'simple' does it? i suppose its a little simpler lol